10 dropsof oleic acid sol. of concentration 1/400cm3 per m3 of alchoal are dropped on a water surface. the circular film thus produced has radius 10 cm .find the molecular size of oleic acid ,if radius of each dropis1.4mm .

Grade:12

1 Answers

Naveen Kumar
9 years ago
The drop has radius,r =1.4mm
volume=4/3*(pi)*r3
and we know the concentration=1/400 cm3/m3
so calculate the number of the molecules of oleic acid in this 10 drops.These molecules have formed the circular surface of area =2*(pi)*10cm
Now we know the radius of the circle and so calculate the area of the circle.
Divide this area by the number of molecules to get the molecular size of the oleic acid.